OpenStudy (anonymous):

6*2^2x - 13*6^x + 6*3^2x = 0 what is the value of X ?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

you sure you typed that right?

myininaya (myininaya):

lets see if we can factor 6u^2-13uv+6v^2=0

OpenStudy (anonymous):

I get x=1

OpenStudy (anonymous):

how

myininaya (myininaya):

see if this helpful to you k?

OpenStudy (anonymous):

so what's your answer ?

myininaya (myininaya):

3u-2v=0 implies we have 3u=2v 2u-3v=0 implies we have 2u=3v remember u=2^x and v=3^x so we have \[3*2^x=2*3^x \] and \[2*2^x=3*3^x\]

myininaya (myininaya):

get your x on one side 3/2=3^x/2^x 3/2=(3/2)^x so x=1 for the first equation

myininaya (myininaya):

2/3=3^x/2^x 2/3=(3/2)^x x=-1 for the second

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but

myininaya (myininaya):

i will leave the checking part to you

OpenStudy (anonymous):

how you 13uv is = -4uv-9uv

myininaya (myininaya):

i was trying to factor to factor you look for factors of a*c that add up to be b 6(6)=36 (-4)(-9)=36 -4-9=-13

OpenStudy (anonymous):

ok